# A path-dependent PDE solver based on signature kernels

### v3

@@ 3. Signature kernels
Our numerical solver is based on so-called signature kernels, a special class of kernels indexed on $\mathrm{BV}_{0}$. In this section we enumerate the main properties of such signature kernels and highlight important aspects related to their numerical evaluation. We begin by recalling definition and properties of classical kernels.
@@ 3.1. Classical kernels
A scalar-valued kernel on a set $\Omega$ is a symmetric function of the form $\kappa:\Omega\times\Omega\to\mathbb{R}$. Such kernel is said to be positive semidefinite if for any $n\in\mathbb{N}$ and points $\omega^{1},...,\omega^{n}\in\Omega$, the Gram matrix $\mathcal{K}:=(\kappa(\omega^{i},\omega^{j}))_{i,j}$ is positive semidefinite, i.e. if for any $\alpha_{1},...,\alpha_{n}\in\mathbb{R}$
@@
Recall that a Hilbert space $\mathcal{H}$ of functions defined on a set $\Omega$ is a reproducing kernel Hilbert space (RKHS) over $\Omega$ if, for each $\omega\in\Omega$, the point evaluation functional at $\omega$, $f\mapsto f(\omega)$, is a continuous linear functional, i.e. there exists a constant $C_{\omega}\geq 0$ such that
@@
(3.1) $\left\langle\kappa(\omega,\cdot),f\right\rangle_{\mathcal{H}}=f(\omega),\quad\text{for all }f\in\mathcal{H}\text{ and }\omega\in\Omega.$
When dealing with questions related to function approximation, one is often interested in understanding how well elements of an RKHS can approximate a function in a given class. One important such class is the family of continuous functions, leading to the notion of cc-universality [53]. If the set $\Omega$ is equipped with an arbitrary topology and $\kappa:\Omega\times\Omega\rightarrow\mathbb{R}$ is a continuous, symmetric, positive semidefinite kernel on $\Omega$ with RKHS $\mathcal{H}$, then we say that $\kappa$ is cc-universal on $\Omega$ if, for every compact subset $\mathcal{K}\subset\Omega$, $\mathcal{H}$ is dense in $\mathcal{C}(\mathcal{K)}$ in the topology of uniform convergence.
Kernels on Euclidean spaces such as linear, polynomial, Gaussian and Matérn kernels have been extensively studied in prior literature. Arguably, the most popular choice (which is also the one we make in our experiments) of cc-universal kernel on +×$\mathbb{R}_{+}\times\mathbb{R}^{d}$ is the radial basis function (RBF) kernels $g_{\sigma}:\mathbb{R}_{+}\times\mathbb{R}^{d}\times\mathbb{R}_{+}\times\mathbb{R}^{d}\to\mathbb{R}$ defined for $(s,x),(t,y)\in\mathbb{R}_{+}\times\mathbb{R}^{d}$ as
@@ 3.2. Signature kernels
(3.4) $\langle a,b\rangle_{T((V))}=\sum_{n=0}^{\infty}\left\langle a_{n},b_{n}\right\rangle_{V^{\otimes n}}$
for any $a=(a_{0},a_{1},...),b=(b_{0},b_{1}...)$ in $T(V)$.
@@
(3.5) $\langle a,b\rangle_{T((V))}=\sum_{n=0}^{\infty}\phi(n)\left\langle a_{n},b_{n}\right\rangle_{V^{\otimes n}}.$
We denote by $\overline{T(V)}$ the Hilbert space obtained by completing $T(V)$ with respect to $\langle\cdot,\cdot\rangle_{T((V))}$. Supposing that $V$ is $d$-dimensional, for some $d>1$, it will be convenient to introduce the set of words $\mathbb{W}$ in the non-commuting letters $\{1,\dots,d\}$, i.e. ordered $m$-tuples $w=(i_{1},i_{2},\cdots,i_{m})\in\{1,\dots,d\}^{m}$ with $|w|:=m\in\mathbb{N}$. The principal ingredient needed to define the signature kernel is a classical transform in stochastic analysis known as the signature. Next we recall its definition. Let $\mathrm{BV}(\mathbb{T},V)$ be defined as in Section 2.1 where the norm in the definition of $p$-variation is the norm in $V$. For any closed interval $[t,t^{\prime}]\subset\mathbb{T}$ and any path $\gamma\in\mathrm{BV}(\mathbb{T},V)$, the signature of $\gamma$ over $[t,t^{\prime}]$ is the following infinite collection
@@
[…]
@@
(3.7) $dY_{s}=Y_{s}\cdot d\gamma_{s}\text{ on $[t,t^{\prime}]$, \ started at }A\in\overline{T(V)}$
@@
$v=\left(v_{0},v_{1},...\right)$ and $w=\left(w_{0},w_{1},...\right)$ in $\overline{T(V)}$ as the element $v\cdot w=\left(z_{0},z_{1},...\right)$ in $\overline{T(V)}$ such that
@@
and where $d\gamma_{t}$ is identified with $(0,d\gamma_{t},0,...)\in\overline{T(V)}$. In particular, the signature $S(\gamma)_{t,t^{\prime}}$ is the unique solution to (3.7) with initial condition $A=\mathbf{1}:=(1,0,0,...)\in\overline{T(V)}$.
@@
For any $a\in V$, the map $f_{a}:\overline{T(V)}\rightarrow L\left(V,\overline{T(V)}\right)$ given by $f_{a}\left(Y\right)=Y\cdot a$ is well defined and linear. Thus, it can easily been seen to satisfy the conditions of the classical versions of Picard-Lindelöf theorem for Young CDEs (e.g. [36, Thm. 1.28]). Therefore, (3.7) admits a unique solution in $\overline{T(V)}$. The fact that any solution $Y_{s}=\left(Y_{s}^{0},Y_{s}^{1},Y_{s}^{2},...\right)$ to (3.7) in $\overline{T(V)}$ must satisfy $Y_{s}^{k}=(A\cdot S\left(\gamma\right)_{t,s})^{k}$ for any $k$ in $\mathbb{N}$ and any $s$ in $\left[t,t^{\prime}\right]$ can be shown by a simple induction on $k$. ∎
@@
[…]
@@
[…]
@@
[46, Theorem 2.5] Let $[s,s^{\prime}],[t,t^{\prime}]\in[0,T]$ and $\gamma\in\mathrm{BV}([t,t^{\prime}],V),\,\tau\in\mathrm{BV}([s,s^{\prime}],V)$ be two paths. Then, the signature kernel of $\gamma,\tau$ is realised as the solution of the following two-parameter integral equation
(3.8) $\kappa_{\text{sig}}(\gamma,\tau)=\kappa_{\text{sig}}(\gamma,\tau)_{t^{\prime},s^{\prime}}\quad\text{where}\quad\kappa_{\text{sig}}(\gamma,\tau)_{r,q}=1+\int_{s}^{q}\int_{t}^{r}\kappa_{\text{sig}}(\gamma,\tau)_{u,v}\langle d\gamma_{u},d\tau_{v}\rangle_{V},$
@@
Several finite difference schemes are available for numerically evaluating solutions to equation (3.8), see [46, Section 3.1] for details.
@@ 3.3. Signature kernels from static kernels
In view of data science applications, instead of taking inner products of signatures directly, it might be beneficial to first lift paths in $V$ to paths with values in the RKHS $\mathcal{H}_{g}$ of a static kernel $g:V\times V\to\mathbb{R}$. More precisely, for any $\gamma\in\mathrm{BV}([t,t^{\prime}],V)$, with $0\leq t<t^{\prime}\leq T$, denote by $\gamma^{g}:[t,t^{\prime}]\to\mathcal{H}_{g}$ the path defined for any $r\in[t,t^{\prime}]$ as $\gamma^{g}_{r}=g(\gamma_{r},\cdot)$. For the signature of a lifted path $\gamma^{g}$ to be well-defined we will require that the latter is at least continuous and of bounded variation, i.e. $\gamma^{g}\in\mathrm{BV}([t,t^{\prime}],\mathcal{H}_{g})$. The following result gives a sufficient condition to ensure this and immediately follows from the definitions of bounded variation and Lipschitz continuity. In particular, in the experimental section, we will make use of the radial basis function kernel as $g$, which is $C^{\infty}$ and Lipschitz continuous.
@@
Let $g:V\times V\to\mathbb{R}$ be a kernel such that for any $v\in V$ the canonical feature map $g(v,\cdot)$ is Lipschitz-continuous and denote by $\mathcal{H}_{g}$ its RKHS. Then, for any $[s,s^{\prime}],[t,t^{\prime}]\subset[0,T]$, the $g$-lifted signature kernel $\kappa^{g}_{\text{sig}}:\mathrm{BV}([t,t^{\prime}],V)\times\mathrm{BV}([s,s^{\prime}],V)\to\mathbb{R}$ is defined as follows
@@
In the sequel, the intervals $[t,t^{\prime}]$ and $[s,s^{\prime}]$ will be understood implicitly from the domain of definition of $\gamma$ and $\tau$. Lemma 3.6 directly yields a kernel trick for the $g$-lifted signature kernel $\kappa^{g}_{\text{sig}}$:
(3.9) $\kappa^{g}_{\text{sig}}(\gamma,\tau)=\kappa^{g}_{\text{sig}}(\gamma,\tau)_{t^{\prime},s^{\prime}}\quad\text{where}\quad\kappa^{g}_{\text{sig}}(\gamma,\tau)_{r,q}=1+\int_{s}^{q}\int_{t}^{r}\kappa^{g}_{\text{sig}}(\gamma,\tau)_{u,v}\left\langle d\gamma^{g}_{u},d\tau^{g}_{v}\right\rangle_{\mathcal{H}_{g}},$
@@ 3.4. Derivatives of static kernels
To approximate solutions to PPDEs, it will be necessary to differentiate a kernel with respect to its input variables. The RBF kernel $g_{\sigma}$, defined in (3.2), is clearly $C^{\infty}$ and its $n^{th}$ order derivative can be obtained using elementary calculus. For example, the first two derivatives admit the following expressions
@@
By the results in [57], for all $x\in V$, the derivatives of $g(x,\cdot)$ are elements of $\mathcal{H}_{g}$. Furthermore, [57, Theorem 1] states that, for any $h\in\mathcal{H}_{g}$ and any multi-index $I$,
@@ 3.5. Derivatives of signature kernels
For all $\gamma,\eta,{\bar{\eta}}\in\mathrm{BV}(\mathbb{T},\mathbb{R}^{e})$, $e\in\mathbb{N}$, the signature and its first two directional derivatives $\partial^{\eta}_{\gamma}S\circ G$ and $\partial^{\eta{\bar{\eta}}}_{\gamma}S\circ G$ exist and are continuous with respect to the $\mathcal{C}_{0}$-norm. Moreover, the derivatives satisfy the CDEs (A.9) and (A.12).
@@ 3.6. Computations of kernel derivatives
$\mathbb{A}_{s,t}(\gamma,\tau)\approx\frac{1}{\Delta s\Delta t}\Big{(}g(\gamma_{s},\tau_{t})-g(\gamma_{s},\tau_{t-\Delta t})-g(\gamma_{s-\Delta s},\tau_{t})+g(\gamma_{s-\Delta s},\tau_{t-\Delta t})\Big{)},$
and similarly for $\mathbb{A}^{\eta}$ and $\mathbb{A}^{\eta,{\bar{\eta}}}.$
@@ 3.7. Universal product kernels
The solutions to our PPDEs are defined on the product space $\Omega=\mathbb{T}\times\mathbb{R}^{d}\times\mathrm{BV}_{0}$. Therefore, we will need to consider product kernels indexed on such product spaces. The following result provides a simple way of constructing kernels that are universal on a product space as products of kernel on the individual factor spaces.
@@
Since the domains $\Omega$ of our PPDEs are of the form $\Omega=\mathbb{T}\times\mathbb{R}^{d}\times\mathrm{BV}_{0}\subset\mathbb{R}^{d+1}\times\mathrm{BV}_{0}$, it would be tempting to consider product kernels of the form $\widetilde{\kappa}^{k,g}:\Omega\times\Omega\to\mathbb{R}$ defined for any $(s,x,\gamma),(t,y,\tau)\in\Omega$ as
$\widetilde{\kappa}^{h,g}((s,x,\gamma),(t,y,\tau)):=k((s,x),(t,y))\kappa^{g}_{\text{sig}}(\gamma,\tau),$
where $k:\mathbb{R}^{d+1}\times\mathbb{R}^{d+1}\to\mathbb{R}$ is some cc-universal kernel on $\mathbb{R}^{d+1}$ such as the RBF kernel in (3.2), and $\kappa^{g}_{\text{sig}}$ is a $g$-lifted signature kernel.
However, signature kernels are only universal when restricted to classes of paths where the signature is an injective map. One such space is $\mathrm{BV}_{0}^{0}\subset\mathrm{BV}_{0}$, the subset of continuous bounded variation paths that are time-augmented and started at the origin 00, endowed with the $1$-variation topology. Denote by $\mathcal{H}_{\kappa^{g}_{\text{sig}}}$ the RKHS associated to the signature kernel $\kappa^{g}_{\text{sig}}$. The next result states that, when restricted to a compact subset $B\subset\mathrm{BV}_{0}^{0}$, elements of $\mathcal{H}_{\kappa}|_{B}$ are dense in $\mathcal{C}(B)$ with the topology of uniform convergence. In other words, $\kappa^{g}_{\text{sig}}$ is cc-universal on $\mathrm{BV}_{0}^{0}$.
@@
The proof of this statement is analogous to [11, Proposition 3.3] with the difference that the point-separation assumption in the Stone-Weirestrass theorem is justified by noting that $S(\gamma)=S(\tau)\iff\gamma=\tau$ when $\gamma,\tau\in\mathrm{BV}_{0}^{0}$
@@
$\kappa^{g,\ell}_{\text{sig}}(\gamma,\tau)=\ell(\gamma_{0},\tau_{0})\kappa^{g}_{\text{sig}}(\gamma-\gamma_{0},\tau-\tau_{0})$
where $\ell:\mathbb{R}^{e}\times\mathbb{R}^{e}\to\mathbb{R}$ is yet another kernel that keeps track of the starting points of the two input paths. The signature kernel $\kappa^{g}_{\text{sig}}$ is cc-universal on $\mathrm{BV}_{0}$ by Lemma 3.17 hence, if $\ell$ is cc-universal, Lemma 3.16 yields that $\kappa^{g,\ell}_{\text{sig}}$ is cc-universal on $\mathrm{BV}_{0}$. Hence, we define the product kernel $\kappa:\Omega\times\Omega\to\mathbb{R}$ for any $(s,x,\gamma),(t,y,\tau)\in\Omega$ as
(3.21) $\kappa((s,x,\gamma),(t,y,\tau))=\kappa^{k,g,\ell}((s,x,\gamma),(t,y,\tau))$ $=k((s,x),(t,y))\kappa^{g,\ell}_{\text{sig}}(\gamma,\tau),$
where $k:\mathbb{R}^{d+1}\times\mathbb{R}^{d+1}\to\mathbb{R}$ is a cc-universal kernel on $\mathbb{R}^{d+1}$. Note that for a path $\check{\gamma}$ which is constant on $[0,s]$ we have $\check{\gamma}-\check{\gamma}_{0}=\check{\gamma}-\check{\gamma}_{s}$ such that $(\check{\gamma}-\gamma_{0})\lvert_{[s,T]}$ is a path starting at zero. Moreover, because the signature is invariant to translation, there exists a measurable map $\tilde{\kappa}^{g}_{s}:\mathrm{BV}_{s}\to\mathbb{R}$ such that for any $\gamma,\tau\in\mathrm{BV}_{0}^{0}$ with $\check{\gamma}$ constant on $[0,s]$ one has
$\kappa^{g}_{\text{sig}}(\gamma,\tau)=\langle S(\gamma^{g})_{0,T},S(\tau^{g})_{0,T}\rangle=\langle S(\gamma^{g})_{0,s}\cdot S(\gamma^{g})_{s,T},S(\tau^{g})_{0,T}\rangle=\tilde{\kappa}^{g}_{s}(\gamma\lvert_{[s,T]},\tau)$
This entails $(s,x,\gamma)\mapsto\kappa((s,x,\gamma),(t,y,\tau))\in\mathcal{C}(\Omega)$, and it is clear what the Gateaux derivative (2.11) means in this context. For future reference and technical reasons, we make the following assumptions on these kernels, satisfied for instance by RBF kernels.

### v4

@@ 3. Signature kernels
Our numerical solver is based on so-called signature kernels, a special class of kernels indexed on $\mathrm{BV}_{0}=\mathcal{C}^{1-var}([0,T],\mathbb{R}^{e})$. In this section we enumerate the main properties of such signature kernels and highlight important aspects related to their numerical evaluation. We begin by recalling definition and properties of classical kernels.
@@ 3.1. Classical kernels
A scalar-valued kernel on a set $\widehat{\Omega}$ is a symmetric function of the form $\kappa:\widehat{\Omega}\times\widehat{\Omega}\to\mathbb{R}$. Such kernel is said to be positive semidefinite if for any $n\in\mathbb{N}$ and points $\omega^{1},...,\omega^{n}\in\widehat{\Omega}$, the Gram matrix $\mathcal{K}:=(\kappa(\omega^{i},\omega^{j}))_{i,j}$ is positive semidefinite, i.e. if for any $\alpha_{1},...,\alpha_{n}\in\mathbb{R}$
@@
Recall that a Hilbert space $\mathcal{H}$ of functions defined on a set $\widehat{\Omega}$ is a reproducing kernel Hilbert space (RKHS) over $\widehat{\Omega}$ if, for each $\omega\in\widehat{\Omega}$, the point evaluation functional at $\omega$, $f\mapsto f(\omega)$, is a continuous linear functional, i.e. there exists a constant $C_{\omega}\geq 0$ such that
@@
(3.1) $\left\langle\kappa(\omega,\cdot),f\right\rangle_{\mathcal{H}}=f(\omega),\quad\text{for all }f\in\mathcal{H}\text{ and }\omega\in\widehat{\Omega}$.
When dealing with questions related to function approximation, one is often interested in understanding how well elements of an RKHS can approximate a function in a given class. One important such class is the family of continuous functions, leading to the notion of cc-universality [55]. If the set $\widehat{\Omega}$ is equipped with an arbitrary topology and $\kappa:\widehat{\Omega}\times\widehat{\Omega}\rightarrow\mathbb{R}$ is a continuous, symmetric, positive semidefinite kernel on $\widehat{\Omega}$ with RKHS $\mathcal{H}$, then we say that $\kappa$ is cc-universal on $\widehat{\Omega}$ if, for every compact subset $\mathcal{X}\subset\widehat{\Omega}$, $\mathcal{H}$ is dense in $\mathcal{C}(\mathcal{X})$ in the topology of uniform convergence.
Kernels on Euclidean spaces such as linear, polynomial, Gaussian and Matérn kernels have been extensively studied in prior literature. Arguably, the most popular choice (which is also the one we make in our experiments) of cc-universal kernel on $\mathbb{R}_{+}\times\mathbb{R}^{d}$ is the radial basis function (RBF) kernels $g_{\sigma}:\mathbb{R}_{+}\times\mathbb{R}^{d}\times\mathbb{R}_{+}\times\mathbb{R}^{d}\to\mathbb{R}$ defined for $(s,x),(t,y)\in\mathbb{R}_{+}\times\mathbb{R}^{d}$ as
@@ 3.2. Signature kernels
$T(V)=\{a\in T((V))\mid a\text{ has a finite number of non-zero entries}\}.$
(3.4) $\langle a,b\rangle_{T(V)}=\sum_{n=0}^{\infty}\left\langle a_{n},b_{n}\right\rangle_{V^{\otimes n}}$
@@
(3.5) $\langle a,b\rangle_{T(V)}=\sum_{n=0}^{\infty}\phi(n)\left\langle a_{n},b_{n}\right\rangle_{V^{\otimes n}}.$
We denote by $\overline{T(V)}$ the Hilbert space obtained by completing $T(V)$ with respect to $\langle\cdot,\cdot\rangle_{T((V))}$. Supposing that $V$ is $d$-dimensional, for some $d>1$, it will be convenient to introduce the set of words $\mathbb{W}$ in the non-commuting letters $\{1,\dots,d\}$, i.e. ordered $m$-tuples $w=(i_{1},i_{2},\cdots,i_{m})\in\{1,\dots,d\}^{m}$ with $|w|:=m\in\mathbb{N}$. The principal ingredient needed to define the signature kernel is a classical transform in stochastic analysis known as the signature. Next we recall its definition. Let $\mathrm{BV}(\mathbb{T},V)$ be defined as in Section 2.1 where the norm in the definition of $p$-variation is the norm in $V$. For any closed interval $t,t^{\prime}]\subset\mathbb{T}$ and any path $\gamma\in\mathrm{BV}(\mathbb{T},V)$, the signature of $\gamma$ over $t,t^{\prime}]$ is the following infinite collection
@@
(3.6) $S(\gamma)^{(w)}_{t,t^{\prime}}=\int_{t<t_{1}<...<t_{m}<t^{\prime}}\,\mathrm{d}\gamma_{t_{1}}^{(i_{1})}...\,\mathrm{d}\gamma_{t_{m}}^{(i_{m})},\quad\text{for any }w=(i_{1},...,i_{m})\in\mathbb{W}.$
@@
(3.7) $dY_{s}=Y_{s}\cdot\,\mathrm{d}\gamma_{s}\text{ on $[t,t^{\prime}]$, \ started at }A\in\overline{T(V)}$
@@
[…]
@@
and where $\mathrm{d}\gamma_{t}$ is identified with $(0,\,\mathrm{d}\gamma_{t},0,...)\in\overline{T(V)}$. In particular, the signature $S(\gamma)_{t,t^{\prime}}$ is the unique solution to (3.7) with initial condition $A=\mathbf{1}:=(1,0,0,...)\in\overline{T(V)}$.
@@
[…]
@@
$S(\gamma)^{(w)}_{t,t^{\prime}}=\int_{t}^{t^{\prime}}S(\gamma)^{(w_{-})}_{t,t^{\prime}}d\gamma^{(i_{m})}_{u},\quad\text{where }w_{-}=(i_{1},...,i_{m-1}).$
@@
Let $0\leq s<s^{\prime}\leq T,\,0\leq t<t^{\prime}\leq T$. Define the signature kernel $\kappa_{\text{sig}}:\mathrm{BV}([t,t^{\prime}],V)\times\mathrm{BV}([s,s^{\prime}],V)\to\mathbb{R}$ as the following symmetric function
@@
[48, Theorem 2.5] Let $s,s^{\prime}],[t,t^{\prime}]\in[0,T]$ and $\gamma\in\mathrm{BV}([t,t^{\prime}],V),\,\tau\in\mathrm{BV}([s,s^{\prime}],V)$ be two paths. Then, the signature kernel of $\gamma,\tau$ is realised as the solution of the following two-parameter integral equation
(3.8) $\kappa_{\text{sig}}(\gamma,\tau)=\kappa_{\text{sig}}(\gamma,\tau)_{t^{\prime},s^{\prime}}\quad\text{where}\quad\kappa_{\text{sig}}(\gamma,\tau)_{r,q}=1+\int_{s}^{q}\int_{t}^{r}\kappa_{\text{sig}}(\gamma,\tau)_{u,v}\langle\,\mathrm{d}\gamma_{u},d\tau_{v}\rangle_{V},$
@@
Several finite difference schemes are available for numerically evaluating solutions to equation (3.8), see [48, Section 3.1] for details.
@@ 3.3. Signature kernels from static kernels
In view of data science applications, instead of taking inner products of signatures directly, it might be beneficial to first lift paths in $V$ to paths with values in the RKHS $\mathcal{H}_{g}$ of a static kernel $g:V\times V\to\mathbb{R}$. More precisely, for any $\gamma\in\mathrm{BV}([t,t^{\prime}],V)$, with $0\leq t<t^{\prime}\leq T$, denote by $\gamma^{g}:[t,t^{\prime}]\to\mathcal{H}_{g}$ the lifted path defined for any $r\in[t,t^{\prime}]$ as $\gamma^{g}_{r}=g(\gamma_{r},\cdot)$. For the signature of a lifted path $\gamma^{g}$ to be well-defined we will require that the latter is at least continuous and of bounded variation, i.e. $\gamma^{g}\in\mathrm{BV}([t,t^{\prime}],\mathcal{H}_{g})$. The following result gives a sufficient condition to ensure this and immediately follows from the definitions of bounded variation and Lipschitz continuity. In particular, in the experimental section, we will make use of the radial basis function kernel as $g$, which is $\mathcal{C}^{\infty}$ and Lipschitz continuous.
@@
Let $g:V\times V\to\mathbb{R}$ be a kernel such that for any $v\in V$ the canonical feature map $g(v,\cdot)$ is Lipschitz-continuous and denote by $\mathcal{H}_{g}$ its RKHS. Then, for any $s,s^{\prime}],[t,t^{\prime}]\subset[0,T]$, the $g$-lifted signature kernel $\kappa^{g}_{\text{sig}}:\mathrm{BV}([t,t^{\prime}],V)\times\mathrm{BV}([s,s^{\prime}],V)\to\mathbb{R}$ is defined as follows
@@
In the sequel, the intervals $t,t^{\prime}]$ and $s,s^{\prime}]$ will be understood implicitly from the domain of definition of $\gamma$ and $\tau$. Lemma 3.6 directly yields a kernel trick for the $g$-lifted signature kernel $\kappa^{g}_{\text{sig}}$:
(3.9) $\kappa^{g}_{\text{sig}}(\gamma,\tau)=\kappa^{g}_{\text{sig}}(\gamma,\tau)_{t^{\prime},s^{\prime}}\quad\text{where}\quad\kappa^{g}_{\text{sig}}(\gamma,\tau)_{r,q}=1+\int_{s}^{q}\int_{t}^{r}\kappa^{g}_{\text{sig}}(\gamma,\tau)_{u,v}\left\langle\,\mathrm{d}\gamma^{g}_{u},d\tau^{g}_{v}\right\rangle_{\mathcal{H}_{g}},$
@@ 3.4. Derivatives of static kernels
To approximate solutions to PPDEs, it will be necessary to differentiate a kernel with respect to its input variables. The RBF kernel $g_{\sigma}$, defined in (3.2), is clearly $\mathcal{C}^{\infty}$ and its $n^{th}$ order derivative can be obtained using elementary calculus. For example, the first two derivatives admit the following expressions
@@
By the results in [60], for all $x\in V$, the derivatives of $g(x,\cdot)$ are elements of $\mathcal{H}_{g}$. Furthermore, [60, Theorem 1] states that, for any $h\in\mathcal{H}_{g}$ and any multi-index $I$,
@@ 3.5. Derivatives of signature kernels
For all $\gamma,\eta,{\bar{\eta}}\in\mathrm{BV}(\mathbb{T},\mathbb{R}^{e})$, $e\in\mathbb{N}$, the signature and its first two directional derivatives $\partial^{\eta}_{\gamma}S\circ G$ and $\partial^{\eta{\bar{\eta}}}_{\gamma}S\circ G$ exist and are continuous with respect to the supremum norm. Moreover, the derivatives satisfy the CDEs (A.9) and (A.12).
@@ 3.6. Computations of kernel derivatives
$\mathbb{A}_{s,t}(\gamma,\tau)$ $=\left\langle\partial_{t}\gamma^{g}_{s},\partial_{t}\tau^{g}_{t}\right\rangle_{\mathcal{H}_{g}}$
$\approx\left\langle\frac{1}{\Delta s}(\gamma^{g}_{s}-\gamma^{g}_{s-\Delta s}),\frac{1}{\Delta t}(\tau^{g}_{t}-\tau^{g}_{t-\Delta t})\right\rangle_{\mathcal{H}_{g}}$
$=\frac{1}{\Delta s\Delta t}\left(\left\langle\gamma^{g}_{s},\tau^{g}_{t}\right\rangle_{\mathcal{H}_{g}}-\left\langle\gamma^{g}_{s},\tau^{g}_{t-\Delta t}\right\rangle_{\mathcal{H}_{g}}\left\langle\gamma^{g}_{s-\Delta s},\tau^{g}_{t}\right\rangle_{\mathcal{H}_{g}}+\left\langle\gamma^{g}_{s-\Delta s},\tau^{g}_{t-\Delta t}\right\rangle_{\mathcal{H}_{g}}\right)$
$:=\frac{1}{\Delta s\Delta t}\Big(\left\langle g(\gamma_{s},\cdot),g(\tau_{t},\cdot)\right\rangle_{\mathcal{H}_{g}}-\left\langle g(\gamma_{s},\cdot),g(\tau_{t-\Delta t},\cdot)\right\rangle_{\mathcal{H}_{g}}$
$\hskip 56.9055pt-\left\langle g(\gamma_{s-\Delta s},\cdot),g(\tau_{t},\cdot)\right\rangle_{\mathcal{H}_{g}}+\left\langle g(\gamma_{s-\Delta s,\cdot)},g(\tau_{t-\Delta t},\cdot)\right\rangle_{\mathcal{H}_{g}}\Big)$
$=\frac{1}{\Delta s\Delta t}\Big(g(\gamma_{s},\tau_{t})-g(\gamma_{s},\tau_{t-\Delta t})-g(\gamma_{s-\Delta s},\tau_{t})+g(\gamma_{s-\Delta s},\tau_{t-\Delta t})\Big),$
where the last equality follows by the reproducing property of the kernel $g$. Similarly for $\mathbb{A}^{\eta}$ and $\mathbb{A}^{\eta,{\bar{\eta}}}$
@@ 3.7. Universal product kernels
The solutions to our PPDEs are defined on the product space $\Omega\subset\mathbb{T}\times\mathbb{R}^{d}\times\mathrm{BV}_{0}$, see (2.2). Therefore, we will need to consider product kernels indexed on such product spaces. The following result provides a simple way of constructing kernels that are universal on a product space as products of kernel on the individual factor spaces.
@@
Since $\mathbb{T}\times\mathbb{R}^{d}\subset\mathbb{R}^{d+1}$, we consider product kernels $\widetilde{\kappa}^{k,g}$ defined for any $(s,x,\gamma),(t,y,\tau)\in\mathbb{T}\times\mathbb{R}^{d}\times\mathrm{BV}_{0}$ as
$\widetilde{\kappa}^{k,g}((s,x,\gamma),(t,y,\tau)):=k((s,x),(t,y))\kappa^{g}_{\text{sig}}(\gamma,\tau),$
where $k:\mathbb{R}^{d+1}\times\mathbb{R}^{d+1}\to\mathbb{R}$ is some cc-universal kernel on $\mathbb{R}^{d+1}$ such as the RBF kernel in (3.2), and $\kappa^{g}_{\text{sig}}$ is a $g$-lifted signature kernel (Definition 3.8).
However, signature kernels are only universal when restricted to classes of paths where the signature is an injective map. One such space is $\mathrm{BV}_{0}^{0}\subset\mathrm{BV}_{0}$, the subset of continuous bounded variation paths that are time-augmented and started at the origin 0, endowed with the $1$-variation topology. Denote by $\mathcal{H}_{\kappa^{g}_{\text{sig}}}$ the RKHS associated to the signature kernel $\kappa^{g}_{\text{sig}}$. The next result states that, when restricted to a compact subset $B\subset\mathrm{BV}_{0}^{0}$, elements of $\mathcal{H}_{\kappa}|_{B}$ are dense in $\mathcal{C}(B)$ with the topology of uniform convergence. In other words, $\kappa^{g}_{\text{sig}}$ is cc-universal on $\mathrm{BV}_{0}^{0}$.
@@
The proof of this statement is analogous to [11, Proposition 3.3] with the difference that the point-separation assumption in the Stone-Weirestrass theorem is justified by noting that $S(\gamma)=S(\tau)\iff\gamma=\tau$ when $\gamma,\tau\in\mathrm{BV}_{0}^{0}$ .
@@
(3.21) $\kappa^{g,\ell}_{\text{sig}}(\gamma,\tau)=\ell(\gamma_{0},\tau_{0})\kappa^{g}_{\text{sig}}(\gamma-\gamma_{0},\tau-\tau_{0})$
where $\ell:\mathbb{R}^{e}\times\mathbb{R}^{e}\to\mathbb{R}$ is yet another kernel that keeps track of the starting points of the two input paths. The signature kernel $\kappa^{g}_{\text{sig}}$ is cc-universal on $\mathrm{BV}_{0}$ by Lemma 3.17 hence, if $\ell$ is cc-universal, Lemma 3.16 yields that $\kappa^{g,\ell}_{\text{sig}}$ is cc-universal on $\mathrm{BV}_{0}$. We thus define the product kernel $\kappa$ for any $(s,x,\gamma),(t,y,\tau)\in\mathbb{T}\times\mathbb{R}^{d}\times\mathrm{BV}_{0}$ as
(3.22) $\kappa((s,x,\gamma),(t,y,\tau))=k((s,x),(t,y))\kappa^{g,\ell}_{\text{sig}}(\gamma,\tau),$
where $k:\mathbb{R}^{d+1}\times\mathbb{R}^{d+1}\to\mathbb{R}$ is a cc-universal kernel on $\mathbb{R}^{d+1}$ and $\kappa^{g,\ell}_{\text{sig}}$ is from (3.21). Note that for a path $\check{\gamma}$ which is constant on $[0,s]$, as we consider in $\Omega$ (2.2), we have $\check{\gamma}-\check{\gamma}_{0}=\check{\gamma}-\check{\gamma}_{s}$ such that $(\check{\gamma}-\gamma_{0})\lvert_{[s,T]}$ is a path starting at zero. Moreover, because the signature is invariant to translation, there exists a measurable map $\tilde{\kappa}^{g}_{s}:\mathrm{BV}_{s}\to\mathbb{R}$ such that for any $\gamma,\tau\in\mathrm{BV}_{0}^{0}$ with $\check{\gamma}$ constant on $[0,s]$ one has
$\kappa^{g}_{\text{sig}}(\gamma,\tau)=\langle S(\gamma^{g})_{0,T},S(\tau^{g})_{0,T}\rangle=\langle S(\gamma^{g})_{0,s}\cdot S(\gamma^{g})_{s,T},S(\tau^{g})_{0,T}\rangle=\tilde{\kappa}^{g}_{s}(\gamma\lvert_{[s,T]},\tau).$
This entails $(s,x,\gamma)\mapsto\kappa((s,x,\gamma),(t,y,\tau))\in\mathcal{C}(\Omega)$, and it is clear what the Gateaux derivative (2.14) means in this context. For future reference and technical reasons, we make the following assumptions on these kernels, satisfied for instance by RBF kernels. Denote by $\mathcal{H}_{\ell},\mathcal{H}_{k}$ the RKHS associated to the kernels $\ell,k$ respectively, and denote by $\varphi:\mathbb{R}^{e}\to\mathcal{H}_{\ell}$ and $\chi:\mathbb{R}^{d+1}\to\mathcal{H}_{k}$ their feature maps, defined as $\varphi(x):=\ell(x,\cdot)$ and $\chi(x):=k(x,\cdot)$.